# Estimating Parameter Fields in Multi-Physics PDEs from Scarce Measurements

### v1

@@ Abstract
Parameterized partial differential equations (PDEs) underpin the mathematical modeling of complex systems in diverse domains, including engineering, healthcare, and physics. A central challenge in using PDEs for real-world applications is to accurately infer the parameters, particularly when the parameters exhibit non-linear and spatiotemporal variations. Existing parameter estimation methods, such as sparse identification and physics-informed neural networks (PINNs), struggle in such cases, especially with nonlinear dynamics, multiphysics interactions, or limited observations of the system response. To address these challenges, we introduce Neptune, a general-purpose method capable of inferring parameter fields from sparse measurements of system responses. Neptune employs independent coordinate neural networks to continuously represent each parameter field in physical space or in state variables. Across various physical and biomedical problems, where direct parameter measurements are prohibitively expensive or unattainable, Neptune significantly outperforms existing methods, achieving robust parameter estimation from as few as 50 observations, reducing parameter estimation errors by two orders of magnitude and dynamic response prediction errors by a factor of ten compared to PINNs. Furthermore, Neptune exhibits superior extrapolation capabilities, enabling accurate predictions in regimes beyond training data where PINN fail. By facilitating reliable and data-efficient parameter inference, Neptune promises broad transformative impacts in engineering, healthcare, and beyond.
@@
Figure 1: Overview of Neptune, a generalizable method for inferring unknown parameter fields from sparse observations of system responses. Physical phenomena are often governed by known PDEs or coupled PDE systems with unknown key parameter fields. Neptune estimates these complex fields using extremely sparse physical response data through a two-stage process to ensure robustness. First, scalar parameters are estimated under an assumed homogeneous field approximation. Second, neural networks (referred to as $\mathcal{N}$) model the local variance of field parameters, along with the previous scalar estimations. Both stages iteratively solve the PDE system numerically, minimizing the error between predicted and observed dynamics via adjoint-based backpropagation. The trained method yields a reliable physical system model capable of predicting phenomena under varying environmental conditions (e.g., source terms, boundary conditions), enabling Extreme Event Resilience, Accurate Extrapolation, and Proactive Strategies as demonstrated in predicting battery thermal runaway behavior, characterizing flow in porous media, and detecting cardiac fibrosis via electrophysiological analysis.
@@
Accurate modeling of physical phenomena is crucial in a wide range of scientific and engineering applications. Differential equations serve as a foundational framework for describing the dynamics of such systems, capturing processes that evolve over time and space. However, the parameters that govern the dynamics of such systems are often unknown or deviate from known values due to factors such as wear and tear, aging, or environmental variations, resulting in significant changes in the system’s behavior. For example, the thermal properties of aged batteries can evolve, increasing the risk of thermal runaway and fires in electric vehicles (EVs). Similarly, in biomedical contexts, alterations in the properties of cardiac tissue may lead to arrhythmias or atrial fibrillation, while variations in the flow properties within porous media can result in groundwater contamination or aquifer depletion. These examples underscore the critical need for parameter estimation, as failing to account for evolving parameters can lead to inaccurate predictions and severe consequences. Beyond accuracy, parameter estimation enables transformative applications, such as simulating future behavior, predicting breaking points, estimating remaining useful life in engineering systems, and creating personalized models for tailored healthcare diagnostics and treatments through more accurate digital twins.
@@
Parameter estimation in multiphysics applications presents several interconnected challenges. The problem is inherently ill-posed, and in many real-world scenarios, direct measurement of parameters is impractical due to cost, inaccessibility, or physical constraints, leaving system responses as the only observable source of information. In addition to these difficulties, available data is often sparse, noisy, or incomplete, limiting the reliability of inference. Moreover, many relevant parameters are spatially distributed and field-dependent rather than scalar, and they are often tightly coupled across scales, making them difficult to disentangle or model independently.
@@
Despite progress in parameter estimation methods, significant limitations remain. Traditional approaches such as finite element updating [1, 2, 3], Bayesian neural networks [4], least squares estimation [5, 6], Kalman filtering [7, 8], Gaussian processes [9, 10, 11], and sparse identification [12, 13] are effective in simplified, single-physics problems. However, they struggle with spatially and temporally varying parameters in nonlinear multiphysics systems. Karhunen–Loève expansions [14, 15, 16], while capable of estimating field parameters, have limited applicability to temporal dynamics [14, 15, 17] and rely on prior assumptions about parameter distributions [16].
In response to these limitations, recent machine learning based approaches such as PINNs [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], Neural Operators [33, 34, 35, 36, 37, 38, 39, 40, 41], and other deep surrogates [42, 43, 44] have gained attention as efficient alternatives for modeling complex, field-dependent parametric systems. However, a growing body of evidence [45] suggests that these methods often rely on weak numerical baselines and suffer from reporting biases, leading to overoptimistic claims of performance and generalization.
PINNs embed governing equations into neural network loss functions, enabling simultaneous learning of system dynamics and parameter fields [46, 47, 48, 49, 20, 50]. While effective in simple cases, they struggle to recover complex parameters from sparse or noisy data, especially in nonlinear, coupled multiphysics settings [51, 17, 52, 53, 54]. PINN-SR [13], which incorporates sparse regression principles similar to SINDy [12] for equation discovery, faces challenges in generalizing to systems with field parameters.
Neural Operators, such as DeepONet [33] and Fourier Neural Operators (FNO) [34], offer an alternative by learning solution operators directly from data and have been increasingly applied to inverse problems [55, 56, 57]. However, most methods first learn the forward map, then perform inverse estimation. Recent approaches such as Neural Inverse Operators, iFNO, and SC-FNO [58, 59, 55, 57] aim to bypass this structure but still require large, high-fidelity datasets [60, 58] with thousands of samples [55, 56]. Moreover, they remain sensitive to data distribution shifts [61] and require extensive hyperparameter tuning in complex problems [40, 61, 60].
@@
Besides PINNs and neural operators, Neural ordinary differential equations (ODEs) [62, 63, 64, 65, 66, 67] model dynamics in the form of differential equations via neural networks. Coupled with adjoint methods [68, 63], they enable memory‑efficient parameter estimation. Yet the majority of existing work targets low‑dimensional ODEs with scalar parameters and known equations, without extending to PDE‑governed systems or spatially/temporally varying fields [69, 70, 71, 72].
@@
Summary of the Neptune results in the context of accuracy for a range of models Problem Err. (N-0%) Err. (N-1%) Err. (N-10%) Description Battery 0.01 ($\pm$0.006) 0.05 ($\pm$0.009) 0.15 ($\pm$0.038) $k(T)$ Porous Media Flow 0.089 ($\pm$0.019) 0.10 ($\pm$0.016) 0.21 ($\pm$0.073) $K(x,y)$ Cardiac Electrophysiology 0.0030 ($\pm$0.00082) 0.016 ($\pm$0.0060) 0.024 ($\pm$0.0052) $D(x,y)$
@@
Table 2: The error is defined as the average absolute error of the estimated parameter compared to the reference. The values in the parentheses denote the noise levels (e.g., noise-free 0%, 1%, and 10%). The samples used in the above cases are around 128. In the battery problem, only the thermal conductivity $k$ estimation is displayed for simplicity. Absolute error is used instead of percentage error, as values near zero can distort the error assessment.
To address these challenges, this article introduces Neptune (Neural Estimation of Parameters in mulTi-physics PDEs Under sparse obSErvations), a robust and generalizable method for field parameter estimation (see Figure 1). Neptune leverages governing equations and integrates numerical solvers with neural network-based parameter inference, achieving both high parameter estimation accuracy and data efficiency, even under extremely sparse and noisy observation settings. More importantly, it employs a two-stage estimation strategy, first recovering coarse scalar parameters, then refining local variations via neural networks, enabling robust and scalable estimation of complex, multiple field parameters in nonlinear multiphysics systems. In addition, Neptune demonstrates strong generalization and extrapolation capabilities, delivering accurate predictions beyond training regions and across diverse scientific domains, all without assuming parameter distributions or full system knowledge. To support this, governing PDEs are discretized into systems of ODEs, enabling efficient adjoint-based optimization through the numerical solver. Meanwhile, the unknown parameter fields are estimated without any prior assumptions about their distributions, further broadening the applicability of Neptune to multiphysics and coupled systems.
@@
In the following sections, we demonstrate the versatility and efficacy of Neptune across various scientific domains, including cardiac electrophysiology, cell migration and proliferation, thermal-chemical battery systems, and flow dynamics in porous media. Through these applications, we highlight the method’s ability to deliver robust, accurate results under data-scarce and noisy conditions, its resilience to data noise, and its adaptability to complex scenarios using real-world datasets. These qualities underscore Neptune’s transformative potential in advancing parameter estimation and predictive modeling for real-world systems. Quantitative results in Tables 1 and 2 further underscore these strengths, illustrating Neptune’s superior performance in parameter estimation across diverse settings.
@@
Figure 2: Parameter estimation in cardiac electrophysiology. a. The training region with Gaussian noise during measurement, and the estimated forward inference of $V$. The predictions are compared between Neptune and PINN. b. The parameter estimation error under varying training data sizes. A dashed line is used as a reference beyond 1,000 data points to indicate that high performance has already been achieved, with no further improvement expected. c. The parameter estimation distribution comparison between Neptune and PINN, for two representative cases of $D$. e. The estimated forward response $V$ comparisons for case 1. When performing temporal extrapolation, PINN predicts values that do not adhere to physics laws.
@@ I Cardiac Electrophysiology
Measurements (with details and comparison to PINN): Additionally, a PINN model [51] is compared to Neptune for joint parameter estimation and dynamic prediction. PINN requires at least 8,000 spatiotemporal samples on a finer $\times100$ mesh, while Neptune achieves comparable results with only 62 sparse observations on a $\times50$ mesh. Figure 2b compares Neptune and PINN in parameter estimation across different training data sizes. Neptune achieves highly accurate and stable parameter estimates with as few as 1,000 samples, while PINN consistently shows higher errors at all data levels. Furthermore, Neptune significantly outperforms PINN even with noisy data, demonstrating superior accuracy under challenging conditions and limited training samples.
@@
Analysis: Figure 2c evaluates the parameter estimation performance of Neptune and PINN across two cases with different parameter distributions. The first column shows the reference distributions, used to generate ground truth data for training. In the second column, Neptune successfully recovers the parameters in both cases. By contrast, in the fourth column, PINN struggles significantly, even with 50 times more data (50,000 vs. 1,000 points), resulting in a mean absolute error (MAE) that is three orders of magnitude larger. In Case 1, PINN performs worse overall, and in Case 2, it fails to capture the random distribution of $D$.
@@
Figure 2d compares state variable $V$ predictions for Case 1 between Neptune and PINN. While PINN performs reasonably well within the training region (first row), its peak error is an order of magnitude higher than Neptune’s. For temporal extrapolation (second row), PINN shows significantly larger errors due to its inaccurate estimation of $D$. As shown in Fig. 2a, Neptune consistently adheres to the underlying physics, providing robust predictions even beyond the training region. It achieves accurate temporal extrapolations beyond 40TUTU, with an MAE of 0.00035, which is over two orders of magnitude lower than PINN’s MAE of 0.066. Moreover, PINN frequently predicts negative $V$ values at multiple timesteps, further indicating its limitations. Additional results for parameter estimation and forward inference under extremely sparse data conditions are detailed in Supplementary Note 2.
@@
Figure 3: Field parameter estimation in thermal runaway problems. a. A prismatic battery, a cylinder battery, and random temperature measurements at the battery surface. Cross-sections AAA-A and BBB-B represent the cross-plane and the in-plane directions, respectively. b. The estimated three field parameters $C_{p},k(y-z)$, and $k(x)$ for the prismatic battery. c. Predicted thermal runaway behaviors with the estimated parameters. The initial temperature of the battery is 25°$C$ (or 298.15KK), and the boundary condition (ambient temperature) is 423.15KK. After accurately estimating the parameters, the model can correctly predict the thermal and chemical responses. d. Evaluation of temperature distribution for the prismatic battery at two typical times as marked in graph c. e. Predicted thermal runaway behaviors for the cylinder battery. f. Evaluation of temperature distribution for the cylinder battery.

### v2

@@ Abstract
Parameterized partial differential equations (PDEs) underpin the mathematical modeling of complex systems in diverse domains, including engineering, healthcare, and physics. A central challenge in using PDEs for real-world applications is to accurately infer the parameters, particularly when the parameters exhibit non-linear and spatiotemporal variations. Existing parameter estimation methods, such as sparse identification, physics-informed neural networks (PINNs), and neural operators, struggle in such cases, especially with nonlinear dynamics, multiphysics interactions, or limited observations of the system response. To address this, we introduce Neptune, a general-purpose method capable of inferring parameter fields from sparse measurements of system responses. Neptune employs independent coordinate neural networks to continuously represent each parameter field in physical space or in state variables. Across various physical and biomedical problems, where direct parameter measurements are prohibitively expensive or unattainable, Neptune significantly outperforms existing methods, achieving robust parameter estimation from as few as 45 measurements, reducing parameter estimation errors by two orders of magnitude and dynamic response prediction errors by a factor of ten to baselines methods. More importantly, it exhibits superior physical extrapolation capabilities, enabling reliable predictions in regimes far beyond the training data. By facilitating data-efficient parameter inference, Neptune promises broad transformative impacts in engineering, healthcare, and beyond.
@@
Figure 1: Overview of Neptune, a generalizable method for inferring unknown parameter fields from sparse observations of system responses. Physical phenomena are often governed by known PDEs or coupled PDE systems with unknown key parameter fields. Neptune estimates these complex fields using extremely sparse physical response data through a two-stage process to ensure robustness. First, scalar parameters are estimated under an assumed homogeneous field approximation. Second, neural networks (referred to as $\mathcal{N}$) model the local variance of parameter fields, along with the previous scalar estimations. Both stages iteratively solve the PDE system numerically, minimizing the error between predicted and observed dynamics via adjoint-based backpropagation. The trained method yields a reliable physical system model capable of predicting phenomena under varying environmental conditions (e.g., source terms, boundary conditions), enabling Extreme Event Resilience, Accurate Extrapolation, and Proactive Strategies as demonstrated in predicting battery thermal runaway behavior, characterizing flow in porous media, and detecting cardiac fibrosis via electrophysiological analysis.
@@
Accurate modeling of physical phenomena is crucial in a wide range of scientific and engineering applications. Differential equations provide a foundational framework for capturing dynamics that evolve over time and space. However, the governing parameters are often unknown or deviate from known values due to factors such as aging, wear, or environmental variation, resulting in significant changes in system behavior. For example, the thermal properties of aged batteries can evolve, increasing the risk of thermal runaway and fires in electric vehicles (EVs). Similarly, in biomedical contexts, changes in cardiac tissue properties can lead to arrhythmias or atrial fibrillation, while variations in porous media flow properties may cause groundwater contamination or aquifer depletion. These examples underscore the critical need for parameter estimation, as failing to account for evolving parameters can lead to inaccurate predictions and severe consequences. Beyond accuracy, parameter estimation enables transformative applications, such as simulating future behavior, predicting failure points, estimating remaining useful life in engineering systems, and creating personalized models for tailored healthcare diagnostics and treatments through high-fidelity digital twins.
@@
Parameter estimation in multiphysics applications presents several interconnected challenges. The problem is ill-posed due to its inverse nature. In many real-world scenarios, direct parameter measurement is impractical due to cost, inaccessibility, or physical constraints, leaving system responses as the only observable data. In addition, available system response data are often sparse, noisy, or incomplete, limiting reliable parameter recovery. More critically, many relevant parameters are not scalar but spatially distributed or field-dependent, and are often tightly coupled across scales, making them difficult to isolate or model independently.
@@
Despite notable progress in parameter estimation methods, significant limitations remain, particularly when addressing nonlinear multiphysics systems or time-dependent behaviors. Traditional approaches such as finite element updating [1, 2, 3], Bayesian neural networks [4], least squares estimation [5, 6], Kalman filtering [7, 8], Gaussian processes [9, 10, 11], and sparse identification [12, 13] are effective in simplified, single-physics problems. However, they struggle with spatially and temporally varying parameters in nonlinear multiphysics systems. Karhunen–Loève expansions [14, 15, 16] can represent parameter fields, but have limited applicability to time-dependent processes [14, 15, 17] and often require prior assumptions about parameter distributions [16].
Recent machine learning based approaches, such as PINNs [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], Neural Operators [33, 34, 35, 36, 37, 38, 39, 40, 41], and other deep surrogates [42, 43, 44] have emerged as efficient alternatives for modeling complex, field-dependent parametric systems. However, growing evidence [45] shows that these methods often rely on weak numerical baselines and suffer from reporting biases, leading to overoptimistic claims of performance and generalization. PINNs embed governing equations into neural network loss functions, enabling simultaneous learning of system dynamics and parameter fields [46, 47, 48, 49, 20, 50]. While effective in simple cases, PINNs struggle to train or recover complex parameters from extremely sparse or noisy data, particularly in nonlinear, coupled multiphysics systems, due to the dual objectives of simultaneous forward and inverse modeling [51, 17, 52, 53, 54]. PINN-SR [13], which incorporates sparse regression principles (SINDy) [12] for parameter and equation discovery, faces challenges in generalizing to systems with parameter fields. Neural Operators, such as DeepONet [33] and Fourier Neural Operators (FNO) [34], learn solution maps directly from data and have been extended to inverse problems [55, 56, 57, 58], yet often require large, high-fidelity datasets comprising thousands of samples [59, 60] and are sensitive to training data distribution [61] and hyperparameter choices [40, 61, 59].
@@
A separate line of work builds on the differential equation framework by embedding neural networks into the governing dynamics, as in neural ODEs [62, 63, 64, 65, 66, 67] and universal differential equations [68]. Coupled with adjoint methods [68, 63], they enable memory‑efficient parameter estimation. Yet the majority of existing work targets low‑dimensional ODEs with scalar parameters and known equations, without extending to PDE‑governed systems or spatially/temporally varying parameter fields [69, 70, 71, 72]. One underlying reason is the high dimensionality and ill‑posedness of field estimation, where directly optimizing all local parameters leads to an enormous search space and unstable convergence. These challenges highlight the need for structured estimation strategies that can first capture dominant system behavior, before resolving more localized variations.
@@
Problem Err. (N-0%) Err. (N-1%) Err. (N-10%) Description Battery 0.01 ($\pm$0.006) 0.05 ($\pm$0.009) 0.15 ($\pm$0.038) $k(T)$ Porous Media Flow 0.089 ($\pm$0.019) 0.10 ($\pm$0.016) 0.21 ($\pm$0.073) $K(x,y)$ Cardiac Electrophysiology 0.0030 ($\pm$0.00082) 0.016 ($\pm$0.0060) 0.024 ($\pm$0.0052) $D(x,y)$
@@
Table 2: Summary of the Neptune results in the context of accuracy for a range of models. The error is defined as the average absolute error of the estimated parameter compared to the reference. The values in the parentheses denote the noise levels (e.g., noise-free 0%, 1%, and 10%). The samples used in the above cases are around 128. In the battery problem, only the thermal conductivity $k$ estimation is displayed for simplicity. Absolute error is used instead of percentage error, as values near zero can distort the error assessment.
To address these limitations, this article introduces Neptune (Neural Estimation of Parameters in mulTi-physics PDEs Under sparse obSErvations), a generalizable and robust method for parameter field estimation in complex systems (see Figure 1). Neptune extends the universal differential equations framework from simple ODEs to multiphysics and coupled PDEs using finite difference discretization, while remaining compatible with standard numerical solvers. Neptune still utilizes adjoints for optimization, but it combines known governing equations with coordinate neural networks to infer unknown parameter fields, and requires no assumptions about their prior distributions. A core feature of Neptune is a structured two-stage estimation strategy: first recovering coarse scalar parameters, then refining spatially-varying fields via neural networks. This enables scalable recovery of multiple, high-resolution parameter fields in coupled systems.
Neptune inherits the flexibility of standard numerical solvers by discretizing governing PDEs into ODE systems via finite differences, enabling direct integration into existing simulation workflows. It does not rely on any prior assumptions about parameter distributions, and remains robust across a wide range of nonlinear, coupled multiphysics systems. By solving directly from the governing equations, Neptune demonstrates strong generalization and extrapolation capabilities, delivering accurate predictions beyond the training region and across diverse physical domains. These properties allow Neptune to achieve accurate and data-efficient parameter estimation, even under extremely sparse and noisy observations. Our key contributions are as follows:
  • Generalizable framework for multiphysics PDEs. Neptune extends universal differential equations to high-dimensional, coupled systems, without requiring prior assumptions on parameter distributions.
  • Two-stage parameter field estimation. A coarse-to-fine strategy enables scalable recovery of complex, spatially-varying parameter fields.
  • Strong generalization from physics priors. Leveraging known governing equations, Neptune extrapolates reliably beyond training regions and under sparse observations.
@@
In the following sections, we demonstrate Neptune’s versatility and efficacy across various scientific domains, including cardiac modeling, thermal-chemical battery systems, porous media flow, and cell migration. These applications showcase Neptune’s robustness under scarce and noisy data, adaptability to complex real-world settings, and strong predictive accuracy. Quantitative results in Tables 1 and 2 highlight Neptune’s consistent performance in parameter estimation across tasks.
@@
Figure 2: Parameter estimation in cardiac electrophysiology. a. The training region with Gaussian noise during measurement, and the estimated forward inference of $V$. The predictions are compared between Neptune and PINN. b. The parameter estimation error under varying training data sizes. c. Comparison of parameter estimation error plots among Neptune, FNO, and PINN for two representative cases of $D$. d. The estimated forward response $V$ comparisons for case 1. When performing temporal extrapolation, PINN predicts values that do not adhere to physics laws.
@@ I Cardiac Electrophysiology
Measurements (with details and comparison to baselines): Neptune demonstrates exceptional efficiency and precision in joint parameter estimation and dynamic prediction. It achieves highly accurate and stable parameter estimates with as few as 1,000 measurements, maintaining robust performance even when utilizing only 62 sparse measurements on a $50\times 50$ mesh. Furthermore, Neptune significantly outperforms alternative approaches under noisy conditions, demonstrating superior accuracy and stability under challenging, data-limited scenarios. In contrast, the PINN model [51] requires a substantially higher data density to achieve functional results, needing at least 8,000 spatiotemporal measurements on a much finer $100\times 100$ mesh. As illustrated in Figure 2b, PINN consistently exhibits higher estimation errors across all data scales. Beyond estimation accuracy, Neptune maintains physical consistency during temporal extrapolation, whereas PINN tends to generate predictions that violate underlying physiological laws (Figure 2d). Finally, the FNO model was validated on a $30\times 30$ grid over 3 total time steps, totaling 2,700 measurements. While FNO provides a standard operator-learning benchmark, Neptune demonstrates a superior ability to recover parameters $D$ from significantly sparser and noisier measurement sets, as evidenced by the error plots in Figure 2c.
@@
Analysis: Figure 2c evaluates the parameter estimation performance of Neptune, PINN, and FNO across two cases with distinct parameter distributions. The first column displays the reference distributions used as ground truth. In the second column, Neptune successfully recovers the parameters in both cases with high fidelity, achieving a Mean Absolute Error (MAE) of $4.12\times 10^{-6}$ in Case 1 and $2.76\times 10^{-4}$ in Case 2. By contrast, the FNO baseline performs reliably only in Case 1 (MAE: $8.14\times 10^{-4}$), where the geometry aligns with its training distribution of rectangles, triangles, and circles. Case 2 reveals a clear sensitivity to out-of-distribution morphologies; while the FNO correctly identifies the spatial position of the target, it erroneously predicts a shape resembling a training circle rather than the actual underlying geometry, resulting in an increased MAE of $1.26\times 10^{-3}$. In the fourth column, PINN struggles significantly despite utilizing 50 times more data (50,000 vs. 1,000 measurements). In Case 1, PINN yields an MAE of $8.81\times 10^{-3}$ which is about three orders of magnitude larger than Neptune and performs poorly overall. In Case 2, it fails to capture the random distribution of $D$ entirely, with the error further degrading to $3.55\times 10^{-2}$.
@@
Figure 2d compares state variable $V$ predictions for Case 1 between Neptune and PINN. We note that FNO is excluded from this comparison because it cannot perform state variable predictions; unlike the physics-informed frameworks, FNO is trained as an end-to-end black-box estimator where the state variable serves as the input to predict the parameter. Furthermore, FNO is strictly constrained to structured grid data representations, whereas Neptune and PINN utilize flexible pointwise measurements.While PINN performs reasonably well within the training region (Fig. 2d, first row), its peak error remains an order of magnitude higher than that of Neptune. For temporal extrapolation (second row), PINN exhibits significantly larger errors due to its inaccurate estimation of $D$. As shown in Fig. 2a, Neptune consistently adheres to the underlying physics, providing robust predictions even beyond the training window. It achieves accurate temporal extrapolations beyond 40 $TU$ with an MAE of 0.00035 which is over two orders of magnitude lower than the PINN MAE of 0.066. Moreover, PINN frequently predicts non-physical negative $V$ values at multiple timesteps, further highlighting its limitations. A comprehensive analysis of these methodological differences and additional results under extremely sparse measurement conditions are provided in Supplementary Note 2 and 2.2.
@@
Figure 3: Parameter field estimation in thermal runaway problems. a. A prismatic battery, a cylinder battery, and random temperature measurements at the battery surface. Cross-sections AAA-A and BBB-B represent the cross-plane and the in-plane directions, respectively. b. The estimated three parameter fields $C_{p},k(y-z)$, and $k(x)$ for the prismatic battery. c. Predicted thermal runaway behaviors with the estimated parameters. The initial temperature of the battery is 25°$C$ (or 298.15KK), and the boundary condition (ambient temperature) is 423.15KK. After accurately estimating the parameters, the model can correctly predict the thermal and chemical responses. d. Evaluation of temperature distribution for the prismatic battery at two typical times as marked in graph c. e. Predicted thermal runaway behaviors for the cylinder battery. f. Evaluation of temperature distribution for the cylinder battery.